# FlowDAS: A Flow-Based Framework for Data Assimilation

### v1

@@ A.2 The Method For Posterior Estimation
Table S.5: Effect of posterior estimation. The RMSE of vorticity estimate from FlowDAS is evaluated on the incompressible Navier-Stokes task using different posterior estimation methods: Euler’s method (1st-order estimation), Heun’s method (2rd-order estimation), as defined in Equation 12, and forecasting without observations (i.e., no correction). For the super-resolution tasks $16^{2}\rightarrow 128^{2}$ and $32^{2}\rightarrow 128^{2}$, both posterior estimation methods significantly outperform forecasting without observations. Among them, the 2nd-order method achieves the lowest RMSE.
Appendix B Experiment Details and Results
@@
B.1 Constructing Training Dataset
Figure S.6: Structure of training data. Consecutive states are paired across multiple simulated trajectories to construct the $\tilde{\bm{I}}_{s}$ and $\bm{R}_{s}$ defined in Equation 13 for training the velocity model $\bm{b}_{s}$.
  1. 1.
  2. 2.
    1. (a)
    2. (b)
  3. 3.
  4. 4.
    1. (a)
      sample states from $\bm{x}_{0:K-1}^{1:T}$ as $\bm{X}_{0}$’s.
    2. (b)
      retrieve the counterpart states from $\bm{x}_{1:K}^{1:T}$ as $\bm{X}_{1}$’s.
    The batches of state pairs $(\bm{X}_{0},\bm{X}_{1})$ are used to construct $\tilde{\bm{I}}_{s}$ and $\bm{R}_{s}$ in Equation 13 for training the velocity model $\bm{b}_{s}$ as described in Algorithm 1.
In summary, we draw many state pairs $(\bm{X}_{0},\bm{X}_{1})$ and estimate the integral over $s$ in Equation 13 by approximated via an empirical expectation over draws of $s\sim U(0,1)$, and for every $s$ and every state pairs $(\bm{X}_{0},\bm{X}_{1})$, we independently compute Equation 14 with $S$ samples of $\bm{z}_{k}$, so the training loss can be rewritten as:
Monte Carlo Sampling Times $J$ Sampling Step Size $\zeta$
Double-well 17 1
Lorenz 1963 21 0.0002
Incompressible Navier Stokes SR 4x 25 1
SR 8x 25 2
SO 5% 25 1
SO 1.5625% 25 1.75
Particle Image Velocimetry 25 1
B.2 Double-well Potential Problem
B.2.1 The Double-well Potential System
@@
B.2.2 Training and Testing Data Generation
@@
B.2.3 Neural Network for Learning the Drift Velocity
@@
B.2.4 Hyperparameters During Inference
@@
B.2.5 Baseline
For this task, we compare our method with SDA and the classic BPF method. For SDA, we fix the window size to two and use the same training data as FlowDAS to ensure fairness. The local score network is implemented as a fully connected neural network, following the architecture proposed in [Bao et al., 2023]. The score network consists of 3 hidden layers, each with a hidden dimension of 50. The model is trained using the AdamW optimizer with a base learning rate of $0.001$, a weight decay of $0.001$, and a linear learning rate scheduler. Training is conducted for $5000$ epochs. For the BPF method, the particle density is set to $16384$.
B.2.6 Results
@@
Compared to the classic BPF method, we find that it struggles to capture dramatic changes in the trajectory. This limitation is primarily due to the small diffusion coefficient $\beta_{d}$, such as $\beta_{d}=0.2$, which causes the predicted filtering density in BPF to concentrate around the mean value dictated by the deterministic part of the transition equation. As a result, extreme cases lying in the tail of the future state distribution $p(\bm{x}_{k+1}\mid\bm{x}_{k})$ are often missed. This phenomenon can be explained by the truncation error arising from the finite particle space [Bao et al., 2023].
@@
B.3 Lorenz 1963
B.3.1 The Lorenz 1963 Dynamic System
@@
Figure S.8: Dynamics modeling evaluation in the Lorenz 1963 task: FlowDAS vs. SDA. FlowDAS produces results that closely match the true states, demonstrating its ability to learn the underlying transition dynamics $p(\bm{x}_{k+1}\mid\bm{x}_{k})$. In contrast, SDA exhibits rapid divergence from the true states. This divergence arises because SDA focus on modeling the joint distribution $p(\bm{x}_{k+1},\bm{x}_{k})$ rather than directly learn the transition dynamics $p(\bm{x}_{k+1}\mid\bm{x}_{k})$, which is inherently less suitable for capturing the system’s underlying dynamics.
B.3.2 Training and Testing Data Generation
We apply the RK4 method described in Section B.3.1 to generate the simulated training and testing data.
B.3.3 Hyperparameters During Inference
@@
B.3.4 Neural Network for Learning the Drift Velocity
@@ B.3.5 Baseline
B.3.6 Accuracy of the Learned Dynamics
@@
B.4 Incompressible Navier-Stokes Flow
B.4.1 Incompressible Navier-Stokes Flow Problem Settings
For the incompressible Navier-Stokes flow problem, we adopted the problem setting from [Chen et al., 2024b] and used the provided training data. The choices of experiment parameters are as follows: $\nu=10^{-3}$, $\alpha=0.1$, $\varepsilon=1$. The stochastic force $\xi$ is defined as:
@@ B.4.1 Incompressible Navier-Stokes Flow Problem Settings
For more details and insights about this problem, see [Chen et al., 2024b].
B.4.2 Hyperparameters During Inference
@@
B.4.3 Neural Network for Learning the Drift Velocity
We used a U-Net architecture to approximate $\bm{b}_{s}$, following the network proposed in [Chen et al., 2024b]. The conditioning on $\bm{X}_{0}$ was implemented through channel concatenation in the input. The architectural details are as follows:
@@ B.4.3 Neural Network for Learning the Drift Velocity
We employ the AdamW optimizer [Loshchilov and Hutter, 2017] with a cosine annealing schedule to reduce the learning rate during training. The base learning rate is set to $2\times 10^{-4}$. Training is conducted with a batch size of $32$ for $10000$ epochs.
B.4.4 Baseline
We compared our method with SDA. For SDA, the score neural network was configured with a temporal window of $2$, an embedding layer dimensionality of $32$, and a base number of feature channels set to $256$. The network depth was $5$, and the activation function used was SiLU [Loshchilov and Hutter, 2017]. The training process used a batch size of 64 and the AdamW optimizer. A linear learning rate scheduler was applied with an initial learning rate of $0.001$, and the weight decay was set to $0.001$. The model was trained for $10000$ epochs.
Figure S.9: The kinetic energy spectrum of: (a) super-resolution task: $16^{2}\rightarrow 128^{2}$; (b) super-resolution task: $32^{2}\rightarrow 128^{2}$; (c) sparse observation task: 1.5625%; (d) sparse observation task: 5%, in the incompressible Navier-Stokes flow task. We present the kinetic energy spectrum of the true state alongside the estimations from FlowDAS and SDA. FlowDAS can produce results that better aligned with the true state in terms of the kinetic energy spectrum, indicating FlowDAS’s superiority in recovering the physics information and effectiveness as a surrogate model for stochastic dynamic systems. In contrast, SDA fails to produce reasonable high-frequency physics, evidenced by the oscillations in the spectrum.
B.4.5 Kinetic Energy Spectrum Analysis
We evaluated the methods from a physics perspective using the kinetic energy spectrum. A better method produces results that align more closely with the kinetic energy spectrum of the true state. The kinetic energy spectrum is computed as follows:
@@ B.4.5 Kinetic Energy Spectrum Analysis
Figure S.10: Data assimilation of incompressible Navier-Stokes flow. Experiments were conducted at observation resolution of $16^{2}$ (left) and observation sparsity level at $1.5625\%$ (right). The goal of super-resolution task is to reconstruct high-resolution vorticity data ($128^{2}$) from low-resolution observations and the goal of sparse observation task is to recover the complete vorticity field.
B.4.6 Additional Results
@@
The kinetic energy spectrum is shown in Figure S.9 (a), and additional visualization results are provided in Figure S.10 (left). FlowDAS effectively reconstructs high-resolution vorticity data from low-resolution observations, while SDA struggles to capture high-frequency physics.
@@
The kinetic energy spectrum is shown in Figure S.9 (b). Similar to the $16^{2}\rightarrow 128^{2}$ task, FlowDAS achieves superior alignment with the true state compared to SDA.
$1.5625\%$ Sparse Observation
The kinetic energy spectrum is presented in Figure S.9 (c), with additional visualization results in Figure S.10 (right). FlowDAS accurately recovers the vorticity field despite the highly sparse observations, significantly outperforming SDA.
@@
The kinetic energy spectrum is shown in Figure S.9 (d). FlowDAS continues to outperform SDA, demonstrating strong recovery of physical information.
@@
B.5 Particle Image Velocimetry
Figure S.11: Reconstructed Vorticity Fields and Error Maps from FlowDAS and SDA. For the example shown in Figure 4, FlowDAS demonstrated superior reconstruction quality compared to the score-based method SDA. Numerically, FlowDAS achieved a lower RMSE of 0.12, outperforming SDA with an RMSE of 0.15.
B.5.1 Experiment Setting
The training data, neural network (including FlowDAS and SDA) is the same as those in the incompressible Navier-stokes flow simulation. Figure 4 shows the standard PIV set up.
B.5.2 Hyperparameters During Inference
Hyperparameters of the inference procedure, specified for the PIV task, are presented in Table S.6.
@@ B.5.2 Hyperparameters During Inference
B.5.3 Additional Results
RMSE
We compared the RMSE of the reconstructed vorticity fields with the true states for the PIV task. Using FlowDAS and the score-based method SDA, we independently generated 10 results and computed the RMSE for each. FlowDAS achieved a lower average RMSE of $0.12$, compared to $0.15$ for SDA. This highlights the ability of FlowDAS in accurately reconstructing vorticity fields for the PIV task.
Error map
Figure S.11 provides the error maps of FlowDAS and SDA for the example shown in Figure 4. In the FlowDAS error map, most regions are darker (indicating lower errors), with fewer bright spots. In contrast, the SDA error map shows more scattered bright areas and generally higher values. These results suggest that FlowDAS recovers the PIV vorticity field more accurately.

### v2

@@ C.2 Posterior Estimation Methods
Table S.7: Effect of posterior estimation. The RMSE of vorticity estimate from FlowDAS is evaluated on the incompressible Navier-Stokes task using different posterior estimation methods as defined in Equations 9 and 10, and forecasting without observations (i.e., no correction). For the super-resolution tasks $16^{2}\rightarrow 128^{2}$ and $32^{2}\rightarrow 128^{2}$, both posterior estimation methods significantly outperform forecasting without observations. Among them, the 2nd-order method achieves the lowest RMSE.
Appendix D Additional Experiment Results
@@
D.1 Double-well Potential Problem
D.1.1 The Double-well Potential System
@@
D.1.2 Training and Testing Data Generation
@@
D.1.3 Neural Network for Learning the Drift Velocity
@@
D.1.4 Hyperparameters During Inference
@@
D.1.5 Baseline
For this task, we compare our method with SDA and the classic BPF method. For SDA, we fix the window size to two and use the same training data as FlowDAS to ensure fairness. The local score network is implemented as a fully connected neural network, following the architecture proposed in [7]. The score network consists of 3 hidden layers, each with a hidden dimension of 50. The model is trained using the AdamW optimizer with a base learning rate of $0.001$, a weight decay of $0.001$, and a linear learning rate scheduler. Training is conducted for $5000$ epochs. For the BPF method, the particle density is set to $16384$.
D.1.6 Results
@@
Compared to the classic BPF method, we find that it struggles to capture dramatic changes in the trajectory. This limitation is primarily due to the small diffusion coefficient $\beta_{d}$, such as $\beta_{d}=0.2$, which causes the predicted filtering density in BPF to concentrate around the mean value dictated by the deterministic part of the transition equation. As a result, extreme cases lying in the tail of the future state distribution $p(\bm{x}_{k+1}\mid\bm{x}_{k})$ are often missed. This phenomenon can be explained by the truncation error arising from the finite particle space [7].
@@
D.2 Lorenz 1963
D.2.1 The Lorenz 1963 Dynamic System
@@
Figure S.7: Forecaseting dynamics modeling evaluation in the Lorenz 1963 task: FlowDAS vs. SDA. FlowDAS produces results that closely match the true states, demonstrating its ability to learn the underlying transition dynamics $p(\bm{x}_{k+1}\mid\bm{x}_{k})$. In contrast, SDA exhibits rapid divergence from the true states. This divergence arises because SDA focus on modeling the joint distribution $p(\bm{x}_{k+1},\bm{x}_{k})$ rather than directly learn the transition dynamics $p(\bm{x}_{k+1}\mid\bm{x}_{k})$, which is inherently less suitable for capturing the system’s underlying dynamics.
D.2.2 Training and Testing Data Generation
We apply the RK4 method described in Section D.2.1 to generate the simulated training and testing data.
D.2.3 Hyperparameters During Inference
@@
D.2.4 Neural Network for Learning the Drift Velocity
@@ D.2.5 Baseline
D.2.6 Accuracy of the Learned Dynamics
@@
D.3 Incompressible Navier-Stokes Flow
D.3.1 Incompressible Navier-Stokes Flow Problem Settings
For the incompressible Navier-Stokes flow problem, we adopted the problem setting from [16] and used the provided training data. The choices of experiment parameters are as follows: $\nu=10^{-3}$, $\alpha=0.1$, $\varepsilon=1$. The stochastic force $\xi$ is defined as:
@@ D.3.1 Incompressible Navier-Stokes Flow Problem Settings
For more details and insights about this problem, see [16].
D.3.2 Hyperparameters During Inference
@@
D.3.3 Neural Network for Learning the Drift Velocity
We use a U-Net architecture to approximate $\bm{b}_{s}$, following the network proposed in [16]. The conditioning on $\bm{X}_{0}$ is implemented through channel concatenation in the input. The architectural details are as follows:
  • Number of states conditioned on: 10.
@@ D.3.3 Neural Network for Learning the Drift Velocity
We employ the AdamW optimizer [35] with a cosine annealing schedule to reduce the learning rate during training. The base learning rate is set to $2\times 10^{-4}$. Training is conducted with a batch size of $32$ for $1000$ epochs.
D.3.4 Baseline
We compare FlowDAS to FNO-DA, Transolver-DA, and SDA.
FNO-DA
The model is configured with $7$ stacked Fourier layers. Each layer processes data in a $128$-dimensional feature space and truncates the Fourier series to include $6$ modes in each spatial direction. A padding of $6$ is applied before the Fourier operations and removed afterward. The model is provided with the previous $10$ states to predict the state at the next time step. We train the model for $1000$ epochs, using a batch size of $64$ and the AdamW optimizer with an initial learning rate of $0.0001$.
Transolver-DA
The model has $8$ layers of transolver blocks whose hidden channels are $128$ and have $32$ slices, with $8$ attention heads. The model is provided with the previous $10$ states to predict the state at the next time step. We train the model for $1000$ epochs using a batch size of $64$, the AdamW optimizer, and a OneCycleLR learning rate scheduler with an initial learning rate of $5e^{-4}$.
SDA
The score neural network is configured with a temporal window of $11$, of which the first $10$ inputs are conditions and the forecast horizon is $1$; an embedding layer dimensionality of $32$; and a base number of feature channels $256$. The network depth is $5$, and the activation function used is SiLU [35]. We train the model using a batch size of $64$ and the AdamW optimizer, with a linear learning rate scheduler (initial learning rate $0.001$) and a weight decay of $0.001$.
During inference, the $\alpha_{1}$ and $\alpha_{2}$ used is varied from task to task for both the FNO-DA and the Transolver-DA. We show the detailed parameter settings in Tables S.9 and S.8. We use an AdamW optimizer for Kalman update with a learning rate of $1e^{-4}$ and a maximum iteration time of $2000$ to ensure the loss converges. Without other indications, we adopt the same optimizer settings for both FNO-DA and Transolver-DA.
$\alpha_{1}$ $\alpha_{2}$
Incompressible Navier Stokes SR 4x 0.63 1
SR 8x 1.59 1
SO 5% 0.10 1
SO 1.5625% 0.29 1
Weather Forecasting 1.22 1
Table S.8: Hyperparameters for FNO-DA in the inference stage of all experiments presented in this study. SR stands for super-resolution task and SO represents sparse observation (inpainting) task
$\alpha_{1}$ $\alpha_{2}$
Incompressible Navier Stokes SR 4x 0.53 1
SR 8x 1.21 1
SO 5% 0.14 1
SO 1.5625% 0.22 1
Weather Forecasting 5.00 1
Table S.9: Hyperparameters for Transolver-DA in the inference stage of all experiments presented in this study. SR stands for super-resolution task and SO represents sparse observation (inpainting) task
D.3.5 Kinetic Energy Spectrum Analysis
We evaluate the methods from a physics perspective using the kinetic energy spectrum. A better method produces results that align more closely with the kinetic energy spectrum of the true state. The kinetic energy spectrum is computed as follows:
@@ D.3.5 Kinetic Energy Spectrum Analysis
Figure S.8: The kinetic energy spectrum of: (a) super-resolution task: $16^{2}\rightarrow 128^{2}$; (b) super-resolution task: $32^{2}\rightarrow 128^{2}$; (c) sparse observation task: 1.5625%; (d) sparse observation task: 5%, in the incompressible Navier-Stokes flow task. We present the kinetic energy spectrum of the true state alongside the estimations from FlowDAS and baseline methods. FlowDAS can produce results that better aligned with the true state in terms of the kinetic energy spectrum, evidenced by the oscillations in the spectrum of baselines, indicating FlowDAS’s superiority in recovering the physics information and effectiveness as a surrogate model for stochastic dynamic systems.
Figure S.9: Data assimilation of incompressible Navier-Stokes flow—Super-Resolution task (8×8$\times$8 ×). Experiments were conducted at observation resolution of $16^{2}$. The goal of super-resolution task is to reconstruct high-resolution vorticity data ($128^{2}$) from low-resolution observations.
Figure S.10: Data assimilation of incompressible Navier-Stokes flow—Super-Resolution task (4×4$\times$4 ×). Experiments were conducted at observation resolution of $32^{2}$. The goal of super-resolution task is to reconstruct high-resolution vorticity data ($128^{2}$) from low-resolution observations.
Figure S.11: Data assimilation of incompressible Navier-Stokes flow—Sparse-Observation task ($5\%$). Experiments were conducted at observation sparsity level at $5\%$. The goal of sparse observation task is to recover the complete vorticity field.
D.3.6 Additional Results
@@
The kinetic energy spectrum is shown in Figure S.8 (a), and additional visualization results are provided in Figure S.9. FlowDAS effectively reconstructs high-resolution vorticity data from low-resolution observations, while SDA struggles to capture high-frequency physics.
@@
The kinetic energy spectrum is shown in Figure S.8 (b), and additional visualization results are provided in Figure S.10. Similar to the $16^{2}\rightarrow 128^{2}$ task, FlowDAS achieves superior alignment with the true state compared to SDA.
@@
The kinetic energy spectrum is presented in Figure S.8 (d), with additional visualization results in Figure S.11. FlowDAS accurately recovers the vorticity field despite the highly sparse observations, significantly outperforming SDA, demonstrating strong recovery of physical information.
@@
D.4 Particle Image Velocimetry
This section presents a realistic application of our method: Particle Image Velocimetry (PIV). PIV is a widely used optical technique for measuring velocity fields in fluids, with many scientific applications in aerodynamics [55, 57, 31], biological flow studies [30, 51, 21], and medical research [13, 54, 37].
D.4.1 Task Description
In a standard PIV setup, as shown in Figure S.12, fluorescent tracer particles are seeded into a fluid flowing through a channel with transparent walls. A laser sheet illuminates the fluid, and particle movements are recorded by a high-speed camera with adjustable temporal resolution. By analyzing the displacement of these tracer particles, the velocity field within the fluid can be determined at sparse locations. Unlike the task in Section 4.2, which involves recovering dense vorticity fields from sparse vorticity observations, PIV introduces a slightly different DA task: recovering dense vorticity fields ($\bm{x}=\bm{\omega}$) from sparse velocity measurements. This observation model is defined by
$y=\mathcal{A}(\bm{v}(\bm{\omega}))+\bm{\eta},$ (S.41)
Figure S.12: Illustration of the real-world Particle Image Velocimetry (PIV) experiment: The flow is seeded with tracer particles illuminated by a laser sheet, and their movements are captured by a camera to derive the sparse velocity field. The goal here is to recover dense vorticity field from the sparse velocity measurement.
D.4.2 Dataset
In this experiment, we use the same fluid dynamics simulation data from the NS experiments described in Section 4.2. However, we convert the vorticity data to velocity fields via Fourier transform to create synthetic PIV datasets [32]. The particle positions in our simulation are randomly initialized and then perturbed according to the simulated flow motion pattern. In these synthetic images, we assume a particle density of 0.03 particles per pixel, a particle diameter of 3 pixels, and a peak intensity of 255 for each particle in grayscale. The images are processed through a standard PIV pipeline to extract particle locations, match corresponding particles across frames, and compute sparse velocity observations. These sparse measurements are then used in DA to reconstruct the full vorticity fields.
D.4.3 Baselines
We compare our method against the SDA solver. Instead of training new scores or stochastic interpolant networks, we directly adopt the networks trained on vorticity data from the incompressible NS flow experiment to evaluate FlowDAS and SDA on the PIV task.
D.4.4 Experiment Setting
The training data, neural network (including FlowDAS and SDA) is the same as those in the incompressible Navier-stokes flow simulation. Figure S.12 shows the standard PIV set up. Hyperparameters of the inference procedure, specified for the PIV task, are presented in Table S.4.
@@ D.4.4 Experiment Setting
Figure S.13: Data assimilation of Particle Image Velocimetry. The vorticity field is visualized in the same way as in Section 4.2. FlowDAS outperforms SDA in terms of reconstruction fidelity and RMSE, recovering more detailed features even when direct observations are not available. These improvements highlight the potential of FlowDAS for real-world applications.
D.4.5 Results
FlowDAS accurately reconstructs the underlying fluid dynamics from observed particle images, producing vorticity estimates with high precision. As shown in Figure S.13, it outperforms the baseline SDA in terms of reconstruction fidelity. Quantitatively, FlowDAS achieves a lower average RMSE of 0.118, compared to 0.154 for SDA. This performance gap highlights the robustness of FlowDAS and its potential utility in fluid dynamics applications.
D.5 Weather Forecasting
D.5.1 Baselines
We compared FlowDAS to FNO-DA, Transolver-DA, and SDA.
FNO-DA
The model is configured with $7$ stacked Fourier layers. Each layer processes data in a $128$-dimensional feature space and truncates the Fourier series to include 64 modes in each spatial direction. A padding of $6$ is applied before the Fourier operations and removed afterward. The model is provided with the previous $6$ states to predict the state at the next time step. We train the model for $1000$ epochs, using a batch size of $200$ and the AdamW optimizer with an initial learning rate of $0.0001$.
Transolver-DA
The model has $8$ layers of transolver blocks whose hidden channels are $128$ and have $32$ slices, with $8$ attention heads. The model is provided with the previous $6$ states to predict the state at the next time step. We train the model for $1000$ epochs using a batch size of $200$, the AdamW optimizer, and a OneCycleLR learning rate scheduler with an initial learning rate of $5e^{-4}$.
SDA
The score neural network is configured with a temporal window of $7$, of which the first $6$ inputs are conditions, and the forecast horizon is $1$; an embedding layer dimensionality of $32$; and a base number of feature channels $256$. The network depth is $5$, and the activation function used is SiLU [35]. We train the model using a batch size of $200$ and the AdamW optimizer, with a linear learning rate scheduler (initial learning rate $0.001$) and a weight decay of $0.001$.
During inference, the $\alpha_{1}$ and $\alpha_{2}$ used is varied from task to task for both the FNO-DA and the Transolver-DA. We show the detailed parameter settings in Tables S.9 and S.8. We use an AdamW optimizer for Kalman update with a learning rate of $1e^{-4}$ and a maximum iteration time of $2000$ to ensure the loss finally becomes stable. Without other indications, we adopt the same optimizer settings for both FNO-DA and Transolver-DA.
D.5.2 Training Details
We experiment with three backbone architectures—U-Net, FNO, and Transolver—to learn the drift term and compare their performance and provide numerical results in Table S.11. The architectural details are as follows:
  • UNet: as shown in Section D.3.3, the only change is the number of states conditioned on. Here we condition on previous $6$ states.
  • FNO: this model is configured with $7$ stacked Fourier layers. Each layer processes data in a $128$-dimensional feature space and truncates the Fourier series to include $64$ modes in each spatial direction. A padding of $6$ is applied before the Fourier operations and removed afterward. The model is provided with the previous $6$ states, the $\bm{X}_{s}$ and the time $s$ as additional channel concatenated to the previous states and the $\bm{X}_{s}$ to predict the drift term.
  • Transolver: this model has $5$ layers of transolver blocks whose hidden channels are $128$ and have $32$ slices, with $8$ attention heads. The sinusoidal timestep embeddings are applied. The inputs are the previous $6$ states, the $\bm{X}_{s}$ and the time $s$ and the output is the drift term.
Backbone of FlowDAS RMSE $\downarrow$ CSI($\tau_{20}$) (0.3) $\uparrow$ CSI($\tau_{40}$) (0.5) $\uparrow$
FNO 0.053$\pm$0.001 0.718$\pm$0.015 0.568$\pm$0.024
Transolver 0.056$\pm$0.002 0.702$\pm$0.018 0.540$\pm$0.028
UNet 0.056$\pm$0.002 0.703$\pm$0.017 0.540$\pm$0.023
Table S.11: Comparison of backbone network architectures of the drift model on the weather forecasting task. All metrics are averaged over multiple runs. Subtle improvements are observed when using a FNO network as the backbone of the drift model.
D.5.3 Results
We present additional visualizations to demonstrate FlowDAS’s ability of capturing unknown system dynamics on this weather forecasting task in Figures S.14, S.15 and S.16. FlowDAS successfully generates accurate estimates of the weather in the future frames in an autoregressive manner. Table 3 provides numerical results of FlowDAS as well as other baseline methods. FlowDAS consistently outperformed DA baselines both in terms of RMSE and CSI metrics.
Figure S.14: Additional result 1: Data assimilation of weather forecasting (Sparse-Observation task). The underlying PDE of the dynamical system is unknown. This experiment was conducted at an observation sparsity level of $10\%$, showing the capacity of long-trajectory forecasting.
Figure S.15: Additional result 2: Data assimilation of weather forecasting (Sparse-Observation task). The underlying PDE of the dynamical system is unknown. This experiment was conducted at an observation sparsity level of $10\%$, showing the capacity of long-trajectory forecasting.
Figure S.16: Additional result 3: Data assimilation of weather forecasting (Sparse-Observation task). The underlying PDE of the dynamical system is unknown. This experiment was conducted at an observation sparsity level of $10\%$, showing the capacity of long-trajectory forecasting.